OpenStudy (anonymous):

set up a triple integral that gives the volume of the solid bounded by the coordinate planes and the plane 3x + 2y -z = 6

OpenStudy (anonymous):

x =0 , y = 0, z = -6 y = 0, z = 0, x = 2 x = 0, z= 0, y =3 \[\int\limits_{-6}^{0}\int\limits_{0}^{3}\int\limits_{0}^{2} 1 dx dy dz\]

OpenStudy (anonymous):

done

OpenStudy (jamesj):

No. That integral is the volume of a box. But the volume you want is some kind of triangular pyramid. So you need to be much more subtle with the limits of integration.